the value of 4200 j kg-1 k-1 in standard form
1
Expert's answer
2020-09-21T08:29:17-0400

In standart form this value will be written as:


"4200J/(kg\cdot K) = 4.2\times 10^3J/(kg\cdot K)"

or

"4.2\space kJ/(kg\cdot K)"

Answer. "4.2\times 10^3J/(kg\cdot K)" or "4.2\space kJ/(kg\cdot K)".